OpenStudy (anonymous):

PLEASE I AM GIVING MEDALS:))) Emily leaves her physics book of mass 1 kg on the top of the table that is inclined at an angle of 20°. What is the acceleration of the book if the coefficient of kinetic friction between the book and the inclined plane is 0.3? 0.4 m/s2 0.5 m/s2 0.6 m/s2 0.7 m/s2

OpenStudy (surry99):

Can you draw a Free body diagram for this problem?

OpenStudy (anonymous):

The book slides because there is a component of its weight acting along the slope (it means the vertical component of the normal force is less than the book's weight). That force, \(wsin(\theta)\), is acting against kinetic friction, \(f_k\), which is upward. So the net force is: \(F_{net}= wsin(\theta) - \mu_kN\) \(ma = mgsin(\theta) - \mu_kmgcos(\theta)\) \(a= g[sin(\theta) - \mu_kcos(\theta)]\) You can see that mass cancels and you can now figure out what acceleration you get with the slope and coefficient of kinetic friction you have.